OpenStudy (anonymous):

A = Pert, where P is the initial investment and r is the interest rate. 1.How long will it take for $2000 to double if it is invested at 13% interest compounded continuously? 2.What rate of interest compounded continuously is needed for an investment of $500 to grow to $900 in 10 years?

OpenStudy (amistre64):

perts double in ln(2)/r

OpenStudy (amistre64):

ln(900/500)/10 is the rate of the other

OpenStudy (anonymous):

can i get some kind of steps?

OpenStudy (amistre64):

A = Pe^rt ; divide off the P A/P = e^rt ; ln the sides ln(A/P) = rt ; and depending on what your looking for, divide of the rate or the time to solve

OpenStudy (amistre64):

I thought the algebra was self evident ....